OpenStudy (anonymous):

What is the standard form equation of the line passing through (-2, 8) with a slope of 2

OpenStudy (jackellyn):

y-y1=m(x-x1) y-8=2(x-(-2)) y-8=2(x+2) y-8=2x+4 y=2x+4+8 When you combine like terms, what do you get @mustangfav2004 ?

OpenStudy (anonymous):

@jackellyn the answer would be -2x+y=12

OpenStudy (jackellyn):

Very good @mustangfav2004 !